Convert an acceleration of 50 km per hour square into metre per second square
1
Expert's answer
2021-07-19T09:51:37-0400
"a=50\ \dfrac{km}{h^2}\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{1\ h^2}{(3600\ s)^2}=0.0038\ \dfrac{m}{s^2}."